# ConstraintBench: Benchmarking LLM Constraint Reasoning on Direct Optimization

### v1

@@ Abstract
Large language models are increasingly applied to operational decision-making where the underlying structure is constrained optimization. Existing benchmarks evaluate whether LLMs can formulate optimization problems as solver code, but leave open a complementary question. Can LLMs directly produce correct solutions to fully specified constrained optimization problems without access to a solver? We introduce ConstraintBench, a benchmark for evaluating LLMs on direct constrained optimization across 10 operations research domains, with all ground-truth solutions verified by the Gurobi solver. Each task presents a natural-language scenario with entities, constraints, and an optimization objective; the model must return a structured solution that a deterministic verifier checks against every constraint and the solver-proven optimum. We evaluate six frontier models on 200 tasks and find that feasibility, not optimality, is the primary bottleneck. The best model achieves only 65.0% constraint satisfaction, yet feasible solutions average 89 to 96% of the Gurobi-optimal objective. No model exceeds 30.5% on joint feasibility and optimality within 0.1% of the solver reference. Per-domain analysis shows large variation in difficulty, with average feasibility spanning from 83.3% in the production mix domain to 0.8% in the crew assignment domain. Further, systematic failure modes include duration constraint misunderstanding, entity hallucination, and a feasibility-optimality decoupling in facility location and vehicle routing where models achieve high feasibility but 0% optimality. ConstraintBench and all evaluation infrastructure will be publicly released.
@@ 1 Introduction
A growing body of work benchmarks LLMs on optimization formulation, translating natural-language problem descriptions into executable solver code that is then run to check whether the resulting objective matches a reference value (NL4Opt [3], IndustryOR [4], Mamo [5], OptMATH [6], ComplexOR [7], OptiBench [8]). These benchmarks answer an important question, whether a model can write correct solver code, but leave a complementary question unaddressed: Can a model directly produce correct decisions for a fully specified optimization problem, without delegating to a solver at inference time. This question tests fundamentally different capabilities, such as constraint reasoning and combinatorial search rather than mathematical modeling and code generation.
@@
    We introduce ConstraintBench, a benchmark for evaluating LLMs on direct constrained optimization across 10 operations research (OR) domains with 200 verified problem solution pairs. The problems focus on evaluating whether models can reason through constraint interactions and produce feasible, near-optimal solutions as structured outputs.
@@ 3.1 Mixed-Integer Programming
For minimization problems, this definition applies directly; for maximization, we invert the sign. In continuous domains, floating-point arithmetic can produce small negative gaps where a model’s solution appears to slightly outperform the solver reference; we treat $|\text{gap}(x)|<\epsilon$ as optimal for tolerance $\epsilon$. We define conditional optimality as $P(|\text{gap}(x)|<\epsilon\mid x\text{ is feasible})$.
@@ 4.3 Evaluation Protocol
Step 3, independent metric calculation. The verifier re-derives the objective value from the model’s raw decision variables using the same mathematical formulation as the Gurobi model. Critiically the model’s self-reported objective value is never trusted.
@@ 5.2 Aggregate Results
Table 3: Aggregate results (feasibility, optimality, and % optimal).
Model Feasibility (%) Optimality (%) % Optimal Errors
@@
The results demonstrate that feasibility, not optimality, is the primary bottleneck. The best-performing model (GPT-5.2-pro) achieves only 65.0% feasibility, meaning over a third of its solutions violate at least one constraint. Yet among feasible solutions, optimality rates are consistently high (95.2% for GPT-5.2-pro, 96.0% for Claude Opus 4.6), indicating that when models do find the feasible region, they tend to find good solutions within it.
This asymmetry showcases a significant capability gap. Models struggle with constraint satisfaction rather than objective optimization. The gap between the high conditional optimality rates ($\sim$}89 to 96%) and the low joint feasibility-and-optimality rates ($\sim$}8.5 to 30.5%) underscores that finding the feasible region is the harder challenge.
However, the high optimality percentages among feasible solutions should be interpreted with care. This pattern is consistent with a well-known property of combinatorial optimization. For many NP-hard problems, polynomial-time algorithms can find solutions within a constant factor of the optimum, but finding the exact optimum remains intractable [16]. Achieving 90 to 95% of the optimal objective is often straightforward once a feasible solution is found, while closing the remaining gap requires navigating subtle trade-offs. Our strict 0.1% threshold for “% Optimal” reflects this. Even at $\sim$}95% average optimality, only 30.5% of all tasks achieve solutions within 0.1% of the Gurobi reference (see Section 5.4 for per-domain analysis of where this gap is concentrated).
It is also notable that model scale does not straightforwardly predict performance. The smaller o4-mini matches Claude Opus 4.5 on feasibility (50.5%) but achieves lower optimality among feasible solutions (89.0% vs. 94.0%), while Claude Opus 4.6 achieves higher conditional optimality (96.0%) than GPT-5.2 (93.8%) despite lower feasibility (49.0% vs. 64.0%). This suggests that constraint reasoning may be a capability axis partially independent of general model capability.
@@ 5.3 Per-Domain Analysis
Model rankings are unstable across domains. o4-mini achieves the highest feasibility on order fulfillment (90.0%) but the lowest on shift scheduling (55.0%); Claude Opus 4.6 achieves 85 to 100% on production mix and facility location but collapses to 10% on bin packing and vehicle routing. No single model dominates across all domains, suggesting that different models have developed different constraint reasoning strengths.
Three domains (facility location, vehicle routing, and crew assignment) achieve 0% joint optimality across all six models, despite facility location achieving 85.0% average feasibility (and two models at 100%). We analyze this feasibility-optimality decoupling in Section 5.4.
@@
Model Prod. Mix Fac. Loc. Job Shop Order Ful. Shift Sch. Portfolio Bin Pack Vehicle Project Crew
@@
GPT-5.2-pro 95.0 100.0 95.0 70.0 85.0 65.0 70.0 65.0 5.0 0.0
GPT-5.2 90.0 85.0 80.0 80.0 80.0 65.0 50.0 80.0 25.0 5.0
Claude Opus 4.6 85.0 90.0 80.0 70.0 70.0 65.0 10.0 10.0 10.0 0.0
Claude Opus 4.5 80.0 90.0 80.0 80.0 65.0 70.0 20.0 10.0 10.0 0.0
o4-mini 85.0 65.0 60.0 90.0 55.0 70.0 5.0 75.0 0.0 0.0
Gemini 3 Pro 65.0 80.0 65.0 85.0 70.0 65.0 5.0 10.0 0.0 0.0
Average 83.3 85.0 76.7 79.2 70.8 66.7 26.7 41.7 8.3 0.8
@@
GPT-5.2-pro 90.0 0.0 70.0 25.0 55.0 25.0 40.0 0.0 0.0 0.0
GPT-5.2 75.0 0.0 40.0 15.0 45.0 10.0 25.0 0.0 0.0 0.0
Claude Opus 4.6 85.0 0.0 40.0 10.0 40.0 25.0 10.0 0.0 5.0 0.0
Claude Opus 4.5 25.0 0.0 35.0 0.0 45.0 0.0 10.0 0.0 10.0 0.0
o4-mini 30.0 0.0 15.0 10.0 30.0 0.0 0.0 0.0 0.0 0.0
Gemini 3 Pro 35.0 0.0 5.0 5.0 40.0 0.0 0.0 0.0 0.0 0.0
Average 56.7 0.0 34.2 10.8 42.5 10.0 14.2 0.0 2.5 0.0
@@ 5.4 Feasibility-Optimality Decoupling
The clearest finding is the decoupling between feasibility and optimality in facility location and vehicle routing. Facility location has the highest average feasibility of any domain (85.0%, with two models at 100%) yet 0% joint optimality across all six models. Vehicle routing shows a similar pattern for several models. GPT-5.2 achieves 80.0% feasibility but 0% optimality.
@@ 5.5 Failure Mode Breakdown
Category 4, budget and capacity threshold violations (facility location, order fulfillment; 105 violations). Facility location violations (79) are dominated by budget exceedances. Models open slightly too many facilities, pushing total fixed costs above the budget constraint. Order fulfillment violations (26) involve minimum order quantities and conditional vendor constraints. These are “near-miss” errors where solutions are structurally sound but violate resource bounds, often by small margins.
@@ 6.1 Limitations
The 0.1% optimality threshold is strict by design, and the absolute optimality numbers are sensitive to threshold choice. Facility location, for instance, has 85.0% average feasibility but 0% joint optimality at 0.1%; at a 5% threshold, it would rise to approximately 27%. Production mix, by contrast, is threshold-insensitive (68% of feasible solutions are within 0.1%). We report threshold sensitivity analysis in Appendix D to enable readers to assess robustness of the headline numbers.
@@ 7 Conclusion
We introduced ConstraintBench, a benchmark for evaluating LLMs on direct constrained optimization across 10 Operations Research domains with all ground-truth solutions verified by the Gurobi Optimizer. Evaluating six frontier models on 200 tasks, we find that constraint satisfaction, not objective optimization, is the primary bottleneck. The best model achieves 65.0% feasibility, and feasible solutions average 89 to 96% of the Gurobi optimum, yet no model exceeds 30.5% on joint feasibility and optimality. Per-domain analysis reveals a difficulty spectrum spanning two orders of magnitude, from near-solved (production mix, 83.3% average feasibility) to near-impossible (crew assignment, 0.8%), with systematic failure modes that range from structural misunderstanding of domain semantics to targeted constraint reasoning errors. The feasibility-optimality decoupling in facility location (85.0% feasibility, 0% optimality, median gap 9.4%) and vehicle routing demonstrate that achieving feasibility and achieving optimality are distinct challenges that current models address unevenly. By providing solver-verified ground truth, constraint-level evaluation, and detailed failure diagnostics, ConstraintBench offers a rigorous measurement of LLM constraint reasoning and a clear target for improvement.
@@ Appendix B Evaluation Hyperparameters
Appendix C Evaluation Hyperparameters by Provider

### v2

# arXiv:2602.22465v2[cs.AI] 27 Feb 2026

@@ Abstract
Large language models are increasingly applied to operational decision-making where the underlying structure is constrained optimization. Existing benchmarks evaluate whether LLMs can formulate optimization problems as solver code, but leave open a complementary question. Can LLMs directly produce correct solutions to fully specified constrained optimization problems without access to a solver? We introduce ConstraintBench, a benchmark for evaluating LLMs on direct constrained optimization across 10 operations research domains, with all ground-truth solutions verified by the Gurobi solver. Each task presents a natural-language scenario with entities, constraints, and an optimization objective; the model must return a structured solution that a deterministic verifier checks against every constraint and the solver-proven optimum. We evaluate six frontier models on 200 tasks and find that feasibility, not optimality, is the primary bottleneck. The best model achieves only 65.0% feasibility, yet feasible solutions average 89 to 96% of the Gurobi-optimal objective. No model exceeds 30.5% on joint feasibility and optimality within 0.1% of the solver reference. Per-domain analysis shows large variation in difficulty, with average feasibility spanning from 85.0% in the facility location domain to 0.8% in the crew assignment domain. Further, systematic failure modes include duration constraint misunderstanding, entity hallucination, and a feasibility-optimality decoupling in facility location and vehicle routing where models achieve high feasibility but 0% optimality. ConstraintBench and all evaluation infrastructure will be publicly released.
@@ 1 Introduction
A growing body of work benchmarks LLMs on optimization formulation, translating natural-language problem descriptions into executable solver code that is then run to check whether the resulting objective matches a reference value (NL4Opt [3], IndustryOR [4], Mamo [5], OptMATH [6], ComplexOR [7], OptiBench [8]). These benchmarks answer an important question, whether a model can write correct solver code, but leave a complementary question unaddressed: Can a model directly produce correct decisions for a fully specified optimization problem, without delegating to a solver at inference time? This question tests fundamentally different capabilities, such as constraint reasoning and combinatorial search rather than mathematical modeling and code generation.
@@
    We introduce ConstraintBench, a benchmark for evaluating LLMs on direct constrained optimization across 10 operations research (OR) domains with 200 verified problem-solution pairs. The problems focus on evaluating whether models can reason through constraint interactions and produce feasible, near-optimal solutions as structured outputs.
@@ 3.1 Mixed-Integer Programming
For minimization problems, this definition applies directly; for maximization, we invert the sign. In continuous domains, floating-point arithmetic can produce small negative gaps where a model’s solution appears to slightly outperform the solver reference; we treat $|\text{gap}(x)|<\epsilon$ as optimal for tolerance $\epsilon$. We define two optimality metrics. For a feasible solution $x$, the objective quality is $(1-|\text{gap}(x)|)\times 100$, representing the percentage of the solver-optimal objective value achieved. The near-optimality rate is $P(|\text{gap}(x)|<\epsilon\mid x\text{ is feasible})$, the fraction of feasible solutions within tolerance $\epsilon$ of the optimum.
@@ 4.3 Evaluation Protocol
Step 3, independent metric calculation. The verifier re-derives the objective value from the model’s raw decision variables using the same mathematical formulation as the Gurobi model. Critically, the model’s self-reported objective value is never trusted.
@@ 5.2 Aggregate Results
Table 3: Aggregate results (feasibility, objective quality, and % optimal).
Model Feasibility (%) Obj. Quality (%) % Optimal Errors
@@
The results demonstrate that feasibility, not optimality, is the primary bottleneck. The best-performing model (GPT-5.2-pro) achieves only 65.0% feasibility, meaning over a third of its solutions violate at least one constraint. Yet among feasible solutions, objective quality is consistently high (95.2% for GPT-5.2-pro, 96.0% for Claude Opus 4.6), indicating that when models do find the feasible region, they tend to find good solutions within it.
This asymmetry reveals a significant capability gap. Models struggle with constraint satisfaction rather than objective optimization. The gap between the high objective quality ($\sim$}89 to 96%) and the low joint near-optimality rates ($\sim$}8.5 to 30.5%) underscores that finding the feasible region is the harder challenge.
However, the high optimality percentages among feasible solutions should be interpreted with care. This pattern is consistent with a well-known property of combinatorial optimization. For many NP-hard problems, polynomial-time algorithms can find solutions within a constant factor of the optimum, but finding the exact optimum remains intractable [14]. Achieving 90 to 95% of the optimal objective is often straightforward once a feasible solution is found, while closing the remaining gap requires navigating subtle trade-offs. Our strict 0.1% threshold for “% Optimal” reflects this. Even at $\sim$}95% average objective quality, only 30.5% of all tasks achieve solutions within 0.1% of the Gurobi reference (see Section 5.4 for per-domain analysis of where this gap is concentrated).
It is also notable that model scale does not straightforwardly predict performance. The smaller o4-mini matches Claude Opus 4.5 on feasibility (50.5%) but achieves lower objective quality among feasible solutions (89.0% vs. 94.0%), while Claude Opus 4.6 achieves higher objective quality (96.0%) than GPT-5.2 (93.8%) despite lower feasibility (49.0% vs. 64.0%). This suggests that constraint reasoning may be a capability axis partially independent of general model capability.
@@ 5.3 Per-Domain Analysis
Model rankings are unstable across domains. o4-mini achieves the highest feasibility on order fulfillment (90.0%) but the lowest on shift scheduling (55.0%); Claude Opus 4.6 achieves 85% to 90% on production mix and facility location but collapses to 10% on bin packing and vehicle routing. No single model dominates across all domains, suggesting that different models have developed different constraint reasoning strengths.
Three domains (facility location, vehicle routing, and crew assignment) achieve 0% joint optimality across all six models, despite facility location achieving 85.0% average feasibility (and one model at 100%). We analyze this feasibility-optimality decoupling in Section 5.4.
@@
Model Fac. Loc. Prod. Mix Order Ful. Job Shop Shift Sch. Portfolio Vehicle Bin Pack Project Crew
@@
GPT-5.2-pro 100.0 95.0 70.0 95.0 85.0 65.0 65.0 70.0 5.0 0.0
GPT-5.2 85.0 90.0 80.0 80.0 80.0 65.0 80.0 50.0 25.0 5.0
Claude Opus 4.6 90.0 85.0 70.0 80.0 70.0 65.0 10.0 10.0 10.0 0.0
Claude Opus 4.5 90.0 80.0 80.0 80.0 65.0 70.0 10.0 20.0 10.0 0.0
o4-mini 65.0 85.0 90.0 60.0 55.0 70.0 75.0 5.0 0.0 0.0
Gemini 3 Pro Preview 80.0 65.0 85.0 65.0 70.0 65.0 10.0 5.0 0.0 0.0
Average 85.0 83.3 79.2 76.7 70.8 66.7 41.7 26.7 8.3 0.8
@@
GPT-5.2-pro 0.0 90.0 25.0 70.0 55.0 25.0 0.0 40.0 0.0 0.0
GPT-5.2 0.0 75.0 15.0 40.0 45.0 10.0 0.0 25.0 0.0 0.0
Claude Opus 4.6 0.0 85.0 10.0 40.0 40.0 25.0 0.0 10.0 5.0 0.0
Claude Opus 4.5 0.0 25.0 0.0 35.0 45.0 0.0 0.0 10.0 10.0 0.0
o4-mini 0.0 30.0 10.0 15.0 30.0 0.0 0.0 0.0 0.0 0.0
Gemini 3 Pro Preview 0.0 35.0 5.0 5.0 40.0 0.0 0.0 0.0 0.0 0.0
Average 0.0 56.7 10.8 34.2 42.5 10.0 0.0 14.2 2.5 0.0
@@ 5.4 Feasibility-Optimality Decoupling
The clearest finding is the decoupling between feasibility and optimality in facility location and vehicle routing. Facility location has the highest average feasibility of any domain (85.0%, with one model at 100%) yet 0% joint optimality across all six models. Vehicle routing shows a similar pattern for several models. GPT-5.2 achieves 80.0% feasibility but 0% optimality.
@@ 5.5 Failure Mode Breakdown
Category 4, budget and capacity threshold violations (facility location, order fulfillment; 105 violations). Facility location violations (79) are dominated by budget exceedances. Models open slightly too many facilities, pushing total fixed costs above the budget constraint. Order fulfillment violations (26) involve minimum order quantities and conditional vendor constraints. These are “near-miss” errors where solutions are structurally sound but violate resource bounds, often by small margins. The remaining 145 violations span miscellaneous constraint types that do not cluster into a single category.
@@ 6.1 Limitations
The 0.1% optimality threshold is strict by design, and the absolute optimality numbers are sensitive to threshold choice. Facility location, for instance, has 85.0% average feasibility but 0% joint optimality at 0.1%; at a 5% threshold, it would rise to approximately 23%. Production mix, by contrast, is threshold-insensitive (68% of feasible solutions are within 0.1%). We report threshold sensitivity analysis in Appendix D to enable readers to assess robustness of the headline numbers.
@@ 7 Conclusion
We introduced ConstraintBench, a benchmark for evaluating LLMs on direct constrained optimization across 10 Operations Research domains with all ground-truth solutions verified by the Gurobi Optimizer. Evaluating six frontier models on 200 tasks, we find that constraint satisfaction, not objective optimization, is the primary bottleneck. The best model achieves 65.0% feasibility, and feasible solutions average 89 to 96% of the Gurobi optimum, yet no model exceeds 30.5% on joint feasibility and optimality. Per-domain analysis reveals a difficulty spectrum spanning two orders of magnitude, from near-solved (facility location, 85.0% average feasibility) to near-impossible (crew assignment, 0.8%), with systematic failure modes that range from structural misunderstanding of domain semantics to targeted constraint reasoning errors. The feasibility-optimality decoupling in facility location (85.0% feasibility, 0% optimality, median gap 9.41%) and vehicle routing demonstrate that achieving feasibility and achieving optimality are distinct challenges that current models address unevenly. By providing solver-verified ground truth, constraint-level evaluation, and detailed failure diagnostics, ConstraintBench offers a rigorous measurement of LLM constraint reasoning and a clear target for improvement.
@@ Appendix B Evaluation Hyperparameters
Appendix C Provider-Specific Evaluation Configuration